# The Low-Degree Hardness of Finding Large Independent Sets in Sparse Random Hypergraphs

### v2

@@ 1.2 The Low-Degree Framework
We must formally define what it means for such a polynomial to find an independent set in a hypergraph. We follow the notation and definitions from [wein2022optimal], in which a rounding procedure is used to produce an independent set of $\mathcal{H}_{2}(n,p)$ based on the output $f(A)$. For brevity, we just state the following definition in this section for ordinary hypergraphs (the multipartite analog can be inferred as it is nearly identical).
@@
$I$ =$=\big{\{}i\in[n]\,:\,f_{i}(A,\omega)\geq 1\big{\}}$
$\tilde{I}$ =$=\big{\{}i\in I\,:\,\forall e\ni i,\,e\not\subseteq I\big{\}}\,,\,\,\text{ and}$
[…]
@@
[…]
  • there is some edge $e$ containing $i$ such that $e\subseteq I$ ($i$ violates the independence constraint).
@@ 1.3 Main Results
  • We construct the hypergraph $H\sim\mathcal{H}(r,n,p)$ on vertex set $[n]\times[r]$ by including each
@@
Let $H=(V_{1}\cup\cdots\cup V_{r},\,E)$ be an $r$-uniform $r$-partite hypergraph for $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. An independent set $I\subseteq V(H)$ is $\bm{\gamma}$-balanced if $|I\cap V_{i}|=\gamma_{i}|I|$ for each $i\in[r]$. We let $\alpha_{\bm{\gamma}}(H)$ denote the size of the largest $\bm{\gamma}$-balanced independent set in $H$.
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. There exists $d^{*}>0$ such that for any $d\geq d^{*}$, there exists $n^{*}>0$ such that for any $n\geq n^{*}$ and $p=d/n^{r-1}$, the hypergraph $H\sim\mathcal{H}(r,n,p)$ satisfies
$(1-\varepsilon)\left(\frac{\log d}{d\,r^{r-1}(r-1)\prod_{i}\gamma_{i}}\right)^{\frac{1}{r-1}}\,\leq\,\frac{\alpha_{\bm{\gamma}}(H)}{rn}\,\leq\,(1+\varepsilon)\left(\frac{\log d}{d\,r^{r-1}(r-1)\prod_{i}\gamma_{i}}\right)^{\frac{1}{r-1}},$
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Additionally, let $i^{\star}=\arg\max_{i}\gamma_{i}$. There exists $d^{*}>0$ such that for any $d\geq d^{*}$, the following hold.
@@
    $k_{j}\leq\gamma_{j}\,r\,n(1-\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
    there is a degree-$1$ polynomial that $(k_{1},\ldots,k_{r},o_{n}(1),\xi,0)$-optimizes the balanced independent set problem in $\mathcal{H}(r,n,p)$ for $p=d/n^{r-1}$.
  • There exist $\eta,\,C_{1},\,C_{2},\,n^{*}>0$ such that the following holds for any $n\geq n^{*}$, $\xi\geq 1$, $1\leq D\leq\dfrac{C_{1}n}{\xi\log n}$, and $\delta\leq\exp\left(-C_{2}\xi\,D\log n\right)$. If
    $k_{j}\,\geq\,\gamma_{j}\,r\,n(1+\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d\,r^{r-1}(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
@@
Setting $\gamma_{i}=1/r$ for each $i$ provides evidence toward Conjecture 2. For arbitrary $\bm{\gamma}$, we observe a statistical–computational gap of a multiplicative factor of $\gamma_{i^{\star}}^{1/(r-1)}$ for low-degree polynomial algorithms.
@@ 1.4 Proof Overview
Let us first discuss the achievability result (Theorem 1.5). It is a well-known fact that sparse random graphs are locally tree-like. In [wein2022optimal], Wein showed that the $s$-neighborhood of a vertex in $\mathcal{H}_{2}(n,d/n)$ can be well approximated by the Poisson Galton-Watson tree, a random graph model for generating trees. In [pal2021community], Pal and Zhu generalized this model to the hypergraph setting, describing the $r$-uniform Poisson Galton-Watson Hypertree $\mathbb{T}_{d}^{GW}$ (see §2.2 for a formal description). We observe that for $p=d/\binom{n-1}{r-1}$, the hypergraph $\mathcal{H}_{r}(n,p)$ is locally $\mathbb{T}_{d}^{GW}$-like, which implies that the performance of a local algorithm on $\mathcal{H}_{r}(n,p)$ is determined, up to the first order, by the expectation of the corresponding algorithm evaluated at the root of $\mathbb{T}_{d}^{GW}$. Our proof now follows in two steps: first, we describe a local algorithm for finding large independent sets in $\mathbb{T}_{d}^{GW}$ (which can be adapted to one on $\mathcal{H}_{r}(n,p)$ by the aforementioned observation); next, we show how a local algorithm on $\mathcal{H}_{r}(n,p)$ can be well-approximated by a low-degree one, completing the proof. For the first step, we adapt a local algorithm of Nie and Verstraëte on the $\Delta$-regular $r$-uniform hypertree $T_{\Delta}^{r}$ [nie2021randomized] to one on $\mathbb{T}_{d}^{GW}$ for an appropriate $\Delta\coloneqq\Delta(d)>0$. The second step follows a similar approach to that of Wein on graphs.
@@
The proof of our impossibility result (Theorem 1.6) falls into a line of work initiated by Gamarnik and Sudan [gamarnik2020low], who studied local algorithms for independent sets of $\mathcal{H}_{2}(n,p)$. The proof of their impossibility result relies on the so-called Overlap Gap Property (OGP). In particular, they show the following: if $I_{1}$ and $I_{2}$ are “large” independent sets, then either $I_{1}\cap I_{2}$ is a “large” independent set or it is “small”, i.e., either $I_{1}$ and $I_{2}$ contain a lot of common vertices or are nearly disjoint. They then show that if a local algorithm succeeds in finding a large independent set, it can be used to construct two independent sets violating the OGP condition. Rahman and Virág [rahman2017local] improved upon their results by considering a more intricate “forbidden” structure involving many independent sets as opposed to just two. Their approach inspired further works in a number of different areas (see for example [gamarnik2014performance, coja2017walksat, chen2019suboptimality]). For low-degree polynomial algorithms, Wein applied this “ensemble” variant of the OGP in the proof of his impossibility result for independent sets in $\mathcal{H}_{2}(n,p)$ [wein2022optimal]. In our proof, we employ the ensemble-OGP in a similar fashion by considering a forbidden structure that involves many independent sets across many correlated random hypergraphs. We show that with high probability this structure does not exist, and any low-degree algorithm that can construct a large independent set can also be used to construct an instance of this structure, leading to a contradiction.
@@
Let us first discuss the statistical threshold (Theorem 1.8). The proof of the upper bound follows by a standard application of the first moment method. For the lower bound, our approach is inspired by that of Frieze in [frieze1990independence] where he proves a high probability bound on the asymptotic size of the maximum independent set in $\mathcal{H}_{2}(n,p)$ (Krivelevich and Sudakov used a similar approach in [krivelevich1998chromatic] for $\mathcal{H}_{r}(n,p)$). Here, we fix a partition $P$ of the vertex set and we say an independent set $I$ is $P$-independent if $I$ contains at most $1$ vertex from each set in $P$. For an appropriate partition $P$, we show that there exists a $\bm{\gamma}$-balanced $P$-independent set of the desired size with high probability. A similar approach was employed by Perkins and the second author in [perkins2024hardness] where they consider the case when $r=2$. For $r\geq 3$, however, certain combinatorial arguments no longer hold, requiring a more extensive analysis. The details are provided in §4.1.
@@
The low-degree algorithm we describe for our achievability result in Theorem 1.9 is inspired by recent work of the first author who considered balanced independent sets of deterministic $r$-partite hypergraphs [dhawan2023balanced]. While their primary interest was on the existence of large balanced independent sets, the proof is constructive and yields a simple degree-$1$ algorithm. The goal is to adapt this algorithm to construct $\bm{\gamma}$-balanced independent sets in $\mathcal{H}(r,n,p)$. We remark that for $r=2$, a similar approach was used by Perkins and the second author [perkins2024hardness], where they describe an algorithm inspired by work of Chakraborty [chakraborti2023extremal]. In their setting, they note that it must be the case that $\gamma_{i}\leq 1/2$ for some $i\in\{1,2\}$. This allows them to apply Chakraborty’s procedure directly. In order for us to apply the first author’s procedure directly, we would require $\gamma_{i}\leq 1/r$ for at least $r-1$ indices $i\in[r]$ (which need not be the case). Therefore, certain modifications are necessary. This modification is described in detail in §4.2 and yields a degree-$1$ polynomial algorithm in order to prove our achievability result.
@@
The proof of our impossibility result follows a similar approach to that of Theorem 1.6. We once again employ the ensemble-OGP and are able to arrive at the desired contradiction. There is an additional layer of complexity due to the structural constraints on the hypergraph and on the independent sets considered (being $r$-partite and $\bm{\gamma}$-balanced, respectively). Therefore, while the overall strategy of the proof follows identically to that of Theorem 1.6, the analysis differs greatly. For $r=2$, Perkins and the second author were able to reduce their argument to the case when $\gamma_{1}=\gamma_{2}=1/2$. This is due to a similar observation stated in the previous paragraph, i.e., one of the $\gamma_{i}$’s is at least $1/2$. For $r\geq 3$, we would require at least $r-1$ of the $\gamma_{i}$’s to be at least $1/r$ (which need not be the case). Therefore, we cannot apply a similar reduction. Furthermore, as in the case of the proof of Theorem 1.8, the combinatorial arguments are more complex for $r\geq 3$ and so the overall analysis is much more involved.
@@ 1.5 Concluding Remarks
In this work, we consider the computational hardness of finding large (balanced) independent sets in random hypergraphs. We focus on low-degree polynomials, a powerful class of algorithms considered to be a useful proxy for computationally efficient algorithms. For the Erdős–Rényi hypergraph $\mathcal{H}_{r}(n,p)$, we determine the threshold for success of such algorithms in constructing independent sets, recovering and generalizing results of Wein [wein2022optimal]. For the Erdős–Rényi $r$-partite hypergraph $\mathcal{H}(r,n,p)$, we prove a matching threshold for balanced independent sets, generalizing results of Perkins and the second author [perkins2024hardness]. In both problems, our results indicate a statistical–computational gap of a multiplicative factor of $r^{-1/(r-1)}$, that is, the low-degree threshold depends on $r$ for $r$-uniform hypergraphs. This is not entirely expected and cannot be easily deduced from the graph case.
A natural extension of our results would be to consider denser hypergraphs, i.e., larger values of the probability parameter $p$. For $\mathcal{H}_{2}(n,p)$ where $p=\Theta(1)$, Matula and Karp’s results [matula1976largest, karp1976probabilistic] indicate a statistical–computational gap of a multiplicative factor of $1/2$ for independent sets. It is not clear how to describe Karp’s algorithm as a low-degree polynomial and so the question of the low-degree threshold in this regime still remains open. Furthermore, to the best of our knowledge, hypergraph versions of these results have not yet appeared in the literature, which begs the following question: is there a statistical–computational gap of a multiplicative factor of $r^{-1/(r-1)}$ for independent sets in $\mathcal{H}_{r}(n,p)$ when $p=\Theta(1)$?
By now, OGP based methods have become standard in proving low-degree thresholds for optimization problems on graphs (see for example [wein2022optimal, huang2022tight, du2023algorithmic, gamarnik2024hardness, perkins2024hardness]). Our work is the first to apply the low-degree framework and the OGP based proof strategy to optimization problems on random hypergraphs. A potential future line of inquiry would be to adapt this strategy to other optimization problems in this more general setting.

### v3

@@ 1.2 The Low-Degree Framework
We must formally define what it means for such a polynomial to find an independent set in a hypergraph. We follow the notation and definitions from [Wei22], in which a rounding procedure is used to produce an independent set of $\mathcal{H}_{2}(n,p)$ based on the output $f(A)$. For brevity, we just state the following definition in this section for ordinary hypergraphs (the multipartite analog can be inferred as it is nearly identical).
@@
$I$ $=\big\{i\in[n]\,:\,f_{i}(A,\omega)\geq 1\big\}$
$\tilde{I}$ $=\big\{i\in I\,:\,\forall e\ni i,\,e\not\subseteq I\big\}\,,\,\,\text{ and}$
$J$ $=\left\{i\in[n]\,:\,f_{i}(A,\omega)\in\left(\frac{1}{2},\,1\right)\right\}\,.$
@@
  • either $f_{i}(A)\in(1/2,\,1)$, or
  • there is some edge $e$ containing $i$ such that $e\subseteq I$ (ii violates the independence constraint).
@@ 1.3 Main Results
  • We construct the hypergraph $H\sim\mathcal{H}(r,n,p)$ on vertex set $n]\times[r]$ by including each
@@
Let $H=(V_{1}\cup\cdots\cup V_{r},\,E)$ be an $r$-uniform $r$-partite hypergraph for $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)\cap\mathbb{Q}$ and $\sum_{i=1}^{r}\gamma_{i}=1$. An independent set $I\subseteq V(H)$ is $\bm{\gamma}$-balanced if $|I\cap V_{i}|=\gamma_{i}|I|$ for each $i\in[r]$. We let $\alpha_{\bm{\gamma}}(H)$ denote the size of the largest $\bm{\gamma}$-balanced independent set in $H$.
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)\cap\mathbb{Q}$ and $\sum_{i=1}^{r}\gamma_{i}=1$. There exists $d^{*}>0$ such that for any $d\geq d^{*}$, there exists $n^{*}>0$ such that for any $n\geq n^{*}$ and $p=d/n^{r-1}$, the hypergraph $H\sim\mathcal{H}(r,n,p)$ satisfies
$(1-\varepsilon)\left(\frac{\log d}{d(r-1)\prod_{i}\gamma_{i}}\right)^{\frac{1}{r-1}}\,\leq\,\frac{\alpha_{\bm{\gamma}}(H)}{n}\,\leq\,(1+\varepsilon)\left(\frac{\log d}{d(r-1)\prod_{i}\gamma_{i}}\right)^{\frac{1}{r-1}},$
@@
Let $\varepsilon>0$ and $r\in\mathbb{N}$ such that $r\geq 2$, and let $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{r})$ be such that $\gamma_{i}\in(0,1)\cap\mathbb{Q}$ and $\sum_{i=1}^{r}\gamma_{i}=1$. Additionally, let $i^{\star}=\arg\max_{i}\gamma_{i}$. There exists $d^{*}>0$ such that for any $d\geq d^{*}$, the following hold.
@@
    $k_{j}\leq\gamma_{j}\,n(1-\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
    there is a degree-$1$ polynomial that $(k_{1},\ldots,k_{r},\exp\left(-\Omega(n)\right),\xi,0)$-optimizes the balanced independent set problem in $\mathcal{H}(r,n,p)$ for $p=d/n^{r-1}$.
  • There exist $\eta,\,C_{1},\,C_{2},\,n^{*}>0$ such that the following holds for any $n\geq n^{*}$, $\xi\geq 1$, $1\leq D\leq\frac{C_{1}n}{\xi\log n}$, and $\delta\leq\exp\left(-C_{2}\xi\,D\log n\right)$. If
    $k_{j}\,\geq\,\gamma_{j}\,n(1+\varepsilon)\left(\frac{\gamma_{i^{\star}}\log d}{d(r-1)\prod_{i=1}^{r}\gamma_{i}}\right)^{1/(r-1)}\quad\text{for }\quad 1\leq j\leq r,$
@@
Setting $\gamma_{i}=1/r$ for each $i$ provides evidence toward Conjecture 2. For arbitrary $\bm{\gamma}$, we observe a statistical–computational gap of a multiplicative factor of $\gamma_{i^{\star}}^{-1/(r-1)}$ for low-degree polynomial algorithms, matching the result of [PW24] for $r=2$.
@@ 1.4 Proof Overview
Let us first discuss the achievability result (Theorem 1.5). It is a well-known fact that sparse random graphs are locally tree-like. In [Wei22], Wein showed that the $s$-neighborhood of a vertex in $\mathcal{H}_{2}(n,d/n)$ can be well approximated by the Poisson Galton-Watson tree, a random graph model for generating trees. In [PZ21], Pal and Zhu generalized this model to the hypergraph setting, describing the $r$-uniform Poisson Galton-Watson Hypertree $\mathbb{T}_{d}^{GW}$ (see §2.2 for a formal description). We observe that for $p=d/\binom{n-1}{r-1}$, the hypergraph $\mathcal{H}_{r}(n,p)$ is locally $\mathbb{T}_{d}^{GW}$-like, which implies that the performance of a local algorithm on $\mathcal{H}_{r}(n,p)$ is determined, up to the first order, by the expectation of the corresponding algorithm evaluated at the root of $\mathbb{T}_{d}^{GW}$. Our proof now follows in two steps: first, we describe a local algorithm for finding large independent sets in $\mathbb{T}_{d}^{GW}$ (which can be adapted to one on $\mathcal{H}_{r}(n,p)$ by the aforementioned observation); next, we show how a local algorithm on $\mathcal{H}_{r}(n,p)$ can be well-approximated by a low-degree one, completing the proof. For the first step, we adapt a local algorithm of Nie and Verstraëte on the $\Delta$-regular $r$-uniform hypertree $T_{\Delta}^{r}$ [NV21] to one on $\mathbb{T}_{d}^{GW}$ for an appropriate $\Delta\coloneqq\Delta(d)>0$. The second step follows a similar approach to that of Wein on graphs.
@@
The proof of our impossibility result (Theorem 1.6) falls into a line of work initiated by Gamarnik and Sudan [GS14], who studied local algorithms for independent sets of $\mathcal{H}_{2}(n,p)$. The proof of their impossibility result relies on the so-called Overlap Gap Property (OGP). In particular, they show the following: if $I_{1}$ and $I_{2}$ are “large” independent sets, then either $I_{1}\cap I_{2}$ is a “large” independent set or it is “small”, i.e., either $I_{1}$ and $I_{2}$ contain a lot of common vertices or are nearly disjoint. They then show that if a local algorithm succeeds in finding a large independent set, it can be used to construct two independent sets violating the OGP condition. Rahman and Virág [RV17] improved upon their results by considering a more intricate “forbidden” structure involving many independent sets as opposed to just two. Their approach inspired further works in a number of different areas (see for example [GS14a, CHH17, Che+19]). The hardness of random optimization problems within the low-degree polynomial framework was first studied in  [GJW20]. Wein applied this “ensemble” variant of the OGP, first introduced in [Che+19], to establish his impossibility result for independent sets in $\mathcal{H}_{2}(n,p)$ [Wei22]. In our proof, we employ the ensemble-OGP in a similar fashion by considering a forbidden structure that involves many independent sets across many correlated random hypergraphs. We show that with high probability this structure does not exist, and any low-degree algorithm that can construct a large independent set can also be used to construct an instance of this structure, leading to a contradiction.
@@
Let us first discuss the statistical threshold (Theorem 1.8). The proof of the upper bound follows by a standard application of the first moment method. For the lower bound, our approach is inspired by that of Frieze in [Fri90] where he proves a high probability bound on the asymptotic size of the maximum independent set in $\mathcal{H}_{2}(n,p)$ (Krivelevich and Sudakov employed a similar approach in [KS98] for $\mathcal{H}_{r}(n,p)$). The strategy follows a second moment argument which shows that the maximum independent set with an additional property has the desired size with positive probability. The additional property ensures this random variable concentrates well, implying the lower bound with high probability through an application of Azuma’s inequality. A similar approach was employed by Perkins and the second author in [PW24] where they consider the case when $r=2$. For $r\geq 3$, however, certain combinatorial arguments no longer hold, requiring a more extensive analysis. The details are provided in §4.1.
@@
The low-degree algorithm we describe for our achievability result in Theorem 1.9 is inspired by recent work of the first author who considered balanced independent sets of deterministic $r$-partite hypergraphs [Dha25]. While the primary interest of that work is on the existence of large balanced independent sets, the proof is constructive and yields a simple degree-$1$ algorithm. The goal is to adapt this algorithm to construct $\bm{\gamma}$-balanced independent sets in $\mathcal{H}(r,n,p)$. We remark that for $r=2$, a similar approach was used by Perkins and the second author [PW24], where they describe an algorithm inspired by work of Chakraborty [Cha23].
@@
The proof of our impossibility result follows a similar approach to that of Theorem 1.6. We once again employ the ensemble-OGP and are able to arrive at the desired contradiction. There is an additional layer of complexity due to the structural constraints on the hypergraph and on the independent sets considered (being $r$-partite and $\bm{\gamma}$-balanced, respectively). Therefore, while the overall strategy of the proof is identical to that of Theorem 1.6, the analysis differs greatly. We remark that for $r=2$, Perkins and the second author employed a similar approach. They were able to reduce their argument to the case when $\gamma_{1}=\gamma_{2}=1/2$, which substantially simplified their arguments. Unfortunately, such a reduction is not always possible for larger uniformities (one can verify that this reduction is possible only when at least r1r-1 of the $\gamma_{i}$’s satisfy $\gamma_{i}\geq 1/r$). As a result, our analysis is much more involved.
@@ 1.5 Concluding Remarks
In this work, we consider the computational hardness of finding large (balanced) independent sets in random hypergraphs. We focus on low-degree polynomials, a powerful class of algorithms considered to be a useful proxy for computationally efficient algorithms. For the Erdős–Rényi hypergraph $\mathcal{H}_{r}(n,p)$, we determine the threshold for success of such algorithms in constructing independent sets, recovering and generalizing results of Wein [Wei22]. For the Erdős–Rényi $r$-partite hypergraph $\mathcal{H}(r,n,p)$, we prove a matching threshold for balanced independent sets, generalizing results of Perkins and the second author [PW24]. In both problems, our results indicate a statistical–computational gap of a multiplicative factor of $r^{1/(r-1)}$, that is, the low-degree threshold depends on $r$ for $r$-uniform hypergraphs. This is not entirely expected and cannot be easily deduced from the graph case.
A natural extension of our results would be to consider denser hypergraphs, i.e., larger values of the probability parameter $p$. For $\mathcal{H}_{2}(n,p)$ where $p=\Theta(1)$, Matula and Karp’s results [Mat76, Kar76, Wei22] indicate a statistical–computational gap of a multiplicative factor of $2$ for independent sets. It is not clear how to describe Karp’s algorithm as a low-degree polynomial and so the question of the low-degree threshold in this regime still remains open. Furthermore, to the best of our knowledge, hypergraph versions of these results have not yet appeared in the literature, which begs the following question: is there a statistical–computational gap of a multiplicative factor of $r^{1/(r-1)}$ for independent sets in $\mathcal{H}_{r}(n,p)$ when $p=\Theta(1)$?
By now, OGP based methods have become standard in proving low-degree thresholds for optimization problems on graphs (see for example [Wei22, HS22, DGH25, GJW24, PW24]). Our work is the among the first to apply the low-degree framework and the OGP based proof strategy to optimization problems on random hypergraphs. A potential future line of inquiry would be to adapt this strategy to other optimization problems in this more general setting.